OpenStudy (anonymous):

Izzy's heartbeat rate was measured on an EKG machine. After conversion, the function produced could be modeled by a sine function and the wave produced a maximum of 8, minimum of -2, and period of π. Which of the following functions could represent Izzy's EKG read-out?

OpenStudy (anonymous):

a. f(x) = 8 sin πx - 2 b. f(x) = 5 sin πx + 3 c. f(x) = 5 sin 2x + 3 d. f(x) = 2 sin 2x + π

OpenStudy (paounn):

maximum of the A function is 8*1-2 = 6, so it can't be. maximum of the D function is 2*1 +pi ~ 5, so it can't be either. Both b and c are viable. By definition of periodic function, if f(x) has period T, f(kx) has period T/k, it means that the period of function b would be 2, and function c would be pi, which is the required answer.

OpenStudy (anonymous):

Okay, thank you for your help

OpenStudy (anonymous):

period is \(\pi\) so go with \(5\sin(2x)+3\)

OpenStudy (anonymous):

the period of \(\sin(2x)\) is \(\frac{2\pi}{2}=\pi\)